Week of 180312

WLCG Operations Call details

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 Geneva time until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Portal
  • Whenever a particular topic needs to be discussed at the operations meeting requiring information from sites or experiments, it is highly recommended to announce it by email to wlcg-operations@cernSPAMNOTNOSPAMPLEASE.ch to make sure that the relevant parties have the time to collect the required information or invite the right people at the meeting.

Best practices for scheduled downtimes

Monday

Attendance:

  • local: Borja (monitoring), Cristian (storage), Gavin (computing), Ivan (ATLAS), Maarten (SCOD + ALICE)
  • remote: Alexander (NLT1), David B (IN2P3-CC), Di (TRUMF), Gareth (RAL), Jean-Roch (CMS), Jens (NDGF), Kyle (OSG), Marcelo (CNAF), Sang-Un (KISTI), Xavier (KIT)

Experiments round table:

  • ATLAS reports ( raw view) -
    • ATLAS Site jamboree & ATLAS S&C technical meeting - last week
    • Rucio workshop the week before
    • Production - 400k / 600k cores utilization
    • Storage - Almost full. Waiting for DAOD absolution camping

  • CMS reports ( raw view) -
    • Pretty good utilization 240k-220k overall.
      • 180k-160k production
      • 70%-30% prod-analysis
    • Incident on firewall of central collector on March 8 (issue understood, alarm now set), lead ramping down to 120k
    • Recurrent issue with /store/unmerged/ deletion to be followed up on
    • Abnormal level of stageout issues, correlated with singularity deployment ; we’ll have to understand better how to make this kind of roll-out smoother.

  • ALICE -
    • Lowish activity on average

  • LHCb reports ( raw view) -
    • Activity
      • HLT farm fully running
      • 2017 data re-stripping ongoing
      • Stripping 29 reprocessing is ongoing
    • Site Issues
      • CNAF: coming back to life, but storage not working since Sunday evening
    • Tier2D
      • Users with UK certificate problems solved by upgrading xrootd server
        • Maarten: xrootd >= 4.8.0 was needed

Sites / Services round table:

  • ASGC: Unexpected power outage in ASGC happened at 13:35~14:05 on 11 March, 2018 due to the generator in data center wasn't working well during power supply interception by power company. Most servers were recovered so far, yet some disk servers are still down due to hardware issue and result in data transfer failures. Claimed an unscheduled downtime from now to 2018-03-14 23:59:00 for service recovery.
  • BNL:
  • CNAF:
    • LHCb
      • performed tests running MC - Passed
      • Performed tests running Stripping productions - Passed
      • Files have started to be replicated, went good till Sunday afternoon
      • There is an issue with LHCb Storm Frontend, Downtime was created, being resolved, should be ok tomorrow
    • CMS
      • also had trouble with communication to Storm, this case was Gridftp issue, had been solved, should be working now.
      • Singularity is showing an issue, last communications show that it is in CMS side.
    • Since all the experiments are working (with minor tuning), from next week on the CNAF report will return to normal, without the need of the experiments situation sheet:

Recovery followup

Service VO Status Expected restart date Readiness GGUS ticket CNAF comment VO comment
Electric power line - OK Production First line in Production with UPS   The second line is working  
Tape buffer ALICE OK   Production 133582 all ALICE services at CNAF are running in the final configuration. Looks OK
Tape buffer ATLAS OK   Production 131742 Ok  
Tape buffer CMS OK   Production 133515 Ok  
Tape buffer LHCb OK Production Production 133673 OK  
Disk ALICE Parity OK   Production 133582 all ALICE services at CNAF are running in the final configuration. Looks OK
Disk ATLAS Parity OK   Production 131742 Ok  
Disk CMS Degraded Parity   Production 133515 raid5 in a few LUNs, raid6 in the others Disks to be replaced
Disk LHCb Parity OK Production Production 133673 raid6 in all LUNs Data fully copied to new storage, Critical Tapes arrived and tested, MC and Stripping Jobs running, Stage and Replication working except for the hiccup described above
Computing farm -   Ready   CMS: configure singularity on only CINECA farm (DONE); ATLAS: FS mounting on farm nodes and restart LSF queues.   All CE's are operational

  • EGI:
  • FNAL: NTR
  • IN2P3: Tomorrow, IN2P3-CC will be in scheduled maintenance for the whole day.
  • JINR: NTR
  • KISTI: NTR
  • KIT: NTR
  • NDGF: NTR
  • NL-T1:
  • NRC-KI:
  • OSG: NTR
  • PIC:
  • RAL: Since Echo (CEPH) storage was enabled for IPv6 there have been problems accessing it via the RAL FTS service. A problem has now been found in the IPv6 configuration on some nodes and we are hopeful this will resolved the FTS-Echo problems imminently. Also: We had to delay an intervention on our Castor storage for updates to be applied to the back-end databases. This is being re-scheduled for Tuesday 27th March. It will affect Castor storage (including Tape access). Access to Echo storage is not affected.
  • TRIUMF: NTR

  • CERN computing services: NTR
  • CERN storage services:
    • There were crashes of EOS-ATLAS and EOS-CMS last week
    • Due to ongoing repack activities, tape recalls may be slow
  • CERN databases:
  • GGUS: NTR
  • Monitoring:
    • Draft reports for the February availability sent around
  • MW Officer:
  • Networks: NTR
  • Security: NTR

AOB:

Edit | Attach | Watch | Print version | History: r19 < r18 < r17 < r16 < r15 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r19 - 2018-03-12 - MaartenLitmaath
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright &© 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback